# VideoMark: A Distortion-Free Robust Watermarking Framework for Video Diffusion Models

### v2

@@ Abstract
This work introduces VideoMark, a distortion-free robust watermarking framework for video diffusion models. As diffusion models excel in generating realistic videos, reliable content attribution is increasingly critical. However, existing video watermarking methods often introduce distortion by altering the initial distribution of diffusion variables and are vulnerable to temporal attacks, such as frame deletion, due to variable video lengths. VideoMark addresses these challenges by employing a pure pseudorandom initialization to embed watermarks, avoiding distortion while ensuring uniform noise distribution in the latent space to preserve generation quality. To enhance robustness, we adopt a frame-wise watermarking strategy with pseudorandom error correction (PRC) codes, using a fixed watermark sequence with randomly selected starting indices for each video. For watermark extraction, we propose a Temporal Matching Module (TMM) that leverages edit distance to align decoded messages with the original watermark sequence, ensuring resilience against temporal attacks. Experimental results show that VideoMark achieves higher decoding accuracy than existing methods while maintaining video quality comparable to watermark-free generation. The watermark remains imperceptible to attackers without the secret key, offering superior invisibility compared to other frameworks. VideoMark provides a practical, training-free solution for content attribution in diffusion-based video generation. Code and data are available at Project Page.
**footnotetext: Equal Contribution.$\dagger$\$$\dagger$\$footnotetext: Corresponding author.
@@ 1 Introduction
In recent years, diffusion models have revolutionized the landscape of AI-generated content, emerging as the state-of-the-art technology for image and video generation [8, 9, 20]. These models can create highly realistic content that is increasingly indistinguishable from human-created media [19]. The rapid advancement in generation quality has created an urgent need to track and attribute AI-generated content, particularly given growing concerns about copyright infringement and potential misuse [1, 27]. To address these challenges, watermarking techniques have emerged as a crucial solution, providing a reliable mechanism for content traceability and authentication in the era of AI-generated media.
Traditional watermarking methods for both images and videos typically operate as post-processing techniques, where watermarks are embedded after content generation [16, 28]. These methods not only require additional computational overhead but also suffer from limited generalization capabilities. Recent research has shifted towards embedding watermarks during the generation process itself. Leveraging the reversibility of DDIM [21] in diffusion models, several methods have achieved notable success in the image domain by manipulating the initial Gaussian noise—such as Tree-Ring [25]—or by embedding messages directly into the noise distribution, as demonstrated by approaches like Gaussian Shading [26] and PRC-Watermark [7].
However, directly adapting image watermarking techniques to the video domain presents unique challenges. First, video DDIM inversion yields lower accuracy than image-based methods. And as a result, methods like VideoShield [10] repeat watermark patterns in initial noise to enhance detection, but these compromise video quality and watermark invisibility. Second, watermark robustness suffers against temporal attacks like frame deletion or reordering, because treating videos as a single entity fails to localize watermarks temporally. Third, variable video lengths pose difficulties for algorithms relying on fixed noise initialization, limiting scalability.
@@
With these needs in mind, we introduce VideoMark, a distortion free and unbiased watermarking framework designed for video diffusion models. To achieve an imperceptible watermark that preserves the original noise characteristics, VideoMark utilizes pseudorandom error correction (PRC) codes [4]. These codes map the watermark bits directly onto the initialized Gaussian noise for every frame. This specific design ensures the watermark integrates seamlessly, thus fulfilling our first design goal. To enable frame-specific watermarking, VideoMark handles each frame’s watermark independently while embedding a sequential logic. Specifically, we generate an extended watermark message sequence. For each video, a random starting position within this master sequence initializes the first frame’s watermark, and subsequent frames derive their watermarks sequentially. This aligns with our second design objective, facilitating both individualized frame watermarking and temporal coherence.
To accurately extract the watermark, we propose a temporal matching module (TMM), which uses edit distance to align the decoded message with the embedded watermark sequence, thereby improving decoding accuracy. Even when videos undergo temporal attacks such as frame deletion, TMM ensures the robustness and reliability of the watermark.
In our experiments, we evaluate the effectiveness of our watermarking framework across different video diffusion models, demonstrating high decoding accuracy, high-quality generated videos, and strong invisibility. Our watermark achieves higher decoding accuracy compared to VideoShield, which is currently the state-of-the-art watermarking approach for video diffusion models. Additionally, our watermark achieves the best video quality on both the objective video evaluation benchmark VBench[11] and subjective assessments, maintaining parity with watermark-free videos. Importantly, our watermark remains undetectable to attackers without the key, ensuring strong imperceptibility compared to other watermarking frameworks.
In summary, the contributions of this work are summarized as follows:
@@ 2.1 Diffusion Models
The reverse process is learned to gradually denoise from $\mathbf{x}T\sim\mathcal{N}(0,\mathbf{I})$ to generate content. While DDPM [8] introduces stochasticity in each denoising step, DDIM [21] provides an approximately invertible deterministic sampling process:
$\mathbf{x}_{t-1}=\sqrt{\alpha_{t-1}}\left(\frac{\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\alpha_{t}}}\right)+\sqrt{1-\alpha_{t-1}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ (2)
@@ 2.2 Pseudorandom Codes (PRC)
A PRC is a coding scheme that maps messages to statistically random-looking codewords. We adopt the construction from Christ and Gunn [4], which provides security based on the hardness of the Learning Parity with Noise (LPN) problem.
@@
Our implementation supports soft decisions on recovered bits, optimized for robust watermarking (see Appendix A.1 for details).
@@ 2.3 Generative Video Watermarking and Threat Model
2. Decoding: $\hat{m}=\mathcal{D}(V)$, where $\mathcal{D}$ extracts the watermark message $\hat{m}$ from the given video $V$.
3. Detection: $\{p,d\}=\mathcal{T}(m,\hat{m})$, where $\mathcal{T}$ compares the original message $m$ with the decoded message $\hat{m}$. This function outputs a p-value $p$ and a boolean decision $d$ indicating whether the distance between $m$ and $\hat{m}$ is significantly smaller than that between $m$ and a random message.
@@ 3 Proposed Method
Figure 2: The overall framework of VideoMark. In the watermark embedding phase, $\epsilon$ is the standard Gaussian noise sampled randomly. In the I2V task, during the watermark extraction phase, we use the first frame of the video as the prompt to predict the initialization noise.
@@ 3.1 Watermark Generation
1:Input: PRC-key $k$, frames $f$, channels $c$, height $h$, width $w$, message $m$, video diffusion model $\mathcal{M}$, variational autoencoder (VAE) decoder $\mathcal{D}$
2:Output: watermarked video $V$
@@
5:for $i\in 1,...,f$ do
6:    Get frame message $m_{i}$ from $M$ starting from position $p+i$
7:    Applying Encode to sample a PRC codeword $c_{i}\in\left\{-1,1\right\}^{c\times h\times w}$ using $m_{i}$
@@
9:    Compute $\hat{\epsilon_{i}}\leftarrow c_{i}\cdot\lvert\epsilon_{i}\rvert$
@@
11:Generate watermarked video $V$ by denoising latent space $\hat{\epsilon}$ with diffusion model $\mathcal{M}$ and decoding it with VAE decoder $\mathcal{D}$
In this section, we introduce the watermark generation process. VideoMark achieves high invisibility and video quality in diffusion-based watermarking by initializing each frame with pseudo-random Gaussian noise using PRC, followed by DDIM denoising[21] and VAE decoding [13]. To enhance diversity and adapt to varying video lengths, we employ an extended message list with a random start index.
Prior watermarking methods (e.g. VideoShield [10]) often repeat identical noise patterns, compromising pseudo-randomness, reducing watermark bit capacity, and degrading invisibility and video quality. VideoMark addresses this by generating frame-specific pseudo-random initializations. For a video with $f$ frames, dimensions $c\times h\times w$ (channels, height, width), and a message bit $m_{i}^{\prime}$ per frame, the process is as follows. For each frame $i\in\{1,\ldots,f\}$, we sample Gaussian noise $\boldsymbol{\epsilon}_{i}\sim\mathcal{N}(0,\mathbf{I})\in\mathbb{R}^{c\times h\times w}$. Using a PRC key $k$, we encode $m_{i}^{\prime}$ to obtain a codeword $\mathbf{c}_{i}=\text{Encode}(k,m_{i}^{\prime})\in\{-1,1\}^{c\times h\times w}$. The watermarked noise is:
@@
$\hat{\boldsymbol{z}}_{i}^{(t-1)}=\sqrt{\alpha_{t-1}}\left(\frac{\hat{\boldsymbol{z}}_{i}^{(t)}-\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}_{\theta}(\hat{\boldsymbol{z}}_{i}^{(t)},t)}{\sqrt{\alpha_{t}}}\right)+\sqrt{1-\alpha_{t-1}}\boldsymbol{\epsilon}_{\theta}(\hat{\boldsymbol{z}}_{i}^{(t)},t),$ (3)
@@
To adapt to videos of varying lengths and increase diversity, we generate an extended message list $M=[m_{1},\ldots,m_{L}]$, where $L>f_{\text{max}}$ and $f_{\text{max}}$ is the maximum supported frame count. For each video, we sample a start index $p\sim\text{Uniform}(0,L-f)$, selecting messages $m_{i}^{\prime}=M[p+i]$ for $i\in\{1,\ldots,f\}$. These are encoded via PRC to produce $\mathbf{c}_{i}$. The random start index ensures diverse initializations across videos, improving security and reducing detectable patterns, while supporting arbitrary video lengths. This frame-wise approach resists temporal and spatial attacks.
The pipeline is shown in Figure 2 and detailed in Algorithm 1.
@@ 3.2 Watermark Extraction
Figure 3: The framework of Temporal Matching Module. ED($\cdot$) denotes the edit distance function. In the Recovered Index list, $I_{1}$ represents the selected starting position $p$.
@@
Decoding Function $\hat{m}=\mathcal{D}(V)$ extracts the embedded message from a watermarked video $V$ with $f$ frames. We first recover the approximate initial noise for each frame using the DDIM inverse process:
$\tilde{\boldsymbol{\epsilon}}_{i}=\mathcal{M}^{-1}(V_{i})\quad\text{for }i\in\{1,...,f\}$ (4)
@@
where the Decode function extracts the message bit encoded in the sign pattern using the PRC key $k$ (details of the PRC algorithm can be found in Appendix A.1), matching the encoding process where $\hat{\boldsymbol{\epsilon}}_{i}=\mathbf{c}_{i}\cdot|\boldsymbol{\epsilon}_{i}|$ and $\mathbf{c}_{i}\in\{-1,1\}^{c\times h\times w}$. The complete decoded sequence is returned as $\hat{m}=[\hat{m}_{1},\ldots,\hat{m}_{f}]$.
Detection Function $\{p,d\}=\mathcal{T}(m,\hat{m})$ determines whether the decoded message $\hat{m}$ contains the watermark message $m$. We compute the edit distance between these sequences, where the cost of insertion, deletion, and replacement operations is 1. To assess statistical significance, we generate $N$ random sequences $\{r^{1},r^{2},...,r^{N}\}$ and compute their edit distances with $\hat{m}$. The p-value is:
@@
where rank is the position of $d_{\text{edit}}(m,\hat{m})$ among all distances. The detection result is $d=\mathbf{1}_{p<\tau}$ with threshold $\tau$. If the p-value is less than $\tau$, there is a watermark with $m$.
@@
$d_{N}(m_{i},\hat{m}_{j})=\begin{cases}0&\text{if }d_{H}(m_{i},\hat{m}_{j})<0.5\\ 2(d_{H}(m_{i},\hat{m}_{j})-0.5)&\text{if }d_{H}(m_{i},\hat{m}_{j})\geq 0.5\end{cases}$ (8)
This transformation maps distances below 0.5 to 0, while linearly scaling distances from 0.5 and above to the range [0,1], enhancing the sensitivity of our detection mechanism.
Recovery Function $m^{\prime}=\mathcal{R}(m,\hat{m})$ is applied when detection succeeds. During the edit distance calculation, we find both the starting index $s$ and the optimal alignment path between $m$ and $\hat{m}$:
$s,\mathcal{P}=\text{argmin}_{i}\{d_{\text{edit}}(m[i:],\hat{m}),\text{Path}(m[i:],\hat{m})\}$ (9)

### v3

@@ Abstract
This work introduces VideoMark, a distortion-free robust watermarking framework for video diffusion models. As diffusion models excel in generating realistic videos, reliable content attribution is increasingly critical. However, existing video watermarking methods often introduce distortion by altering the initial distribution of diffusion variables and are vulnerable to temporal attacks, such as frame deletion, due to variable video lengths. VideoMark addresses these challenges by employing a pure pseudorandom initialization to embed watermarks, avoiding distortion while ensuring uniform noise distribution in the latent space to preserve generation quality. To enhance robustness, we adopt a frame-wise watermarking strategy with pseudorandom error correction (PRC) codes, using a fixed watermark sequence with randomly selected starting indices for each video. For watermark extraction, we propose a Temporal Matching Module (TMM) that leverages edit distance to align decoded messages with the original watermark sequence, ensuring resilience against temporal attacks. Experimental results show that VideoMark achieves higher decoding accuracy than existing methods while maintaining video quality comparable to watermark-free generation. The watermark remains imperceptible to attackers without the secret key, offering superior invisibility compared to other frameworks. VideoMark provides a practical, training-free solution for content attribution in diffusion-based video generation. Our code and data are available at VideoMark.
@@ 1 Introduction
In recent years, diffusion models have revolutionized the landscape of AI-generated content, emerging as the state-of-the-art technology for image and video generation [ho2020denoising, ho2022video, sohl2015deep, liu2025javisgpt]. These models can create highly realistic content that is increasingly indistinguishable from human-created media [rombach2022high]. The rapid advancement in generation quality has created an urgent need to track and attribute AI-generated content, particularly given growing concerns about copyright infringement and potential misuse [almutairi2022review, zhang2025cohemark]. To address these challenges, watermarking techniques have emerged as a crucial solution for ensuring content traceability and authentication in the era of AI-generated media.
@@
Traditional watermarking methods for both images and videos typically operate as post-processing techniques, where watermarks are embedded after content generation [luo2023dvmark, zhang2019robust, ye5194026cmat]. These methods not only require additional computational overhead but also suffer from limited generalization capabilities. Recent research has shifted towards embedding watermarks during the generation process itself. Leveraging the reversibility of DDIM [song2020denoising], several methods have achieved success in the image domain by manipulating the initial Gaussian noise—e.g., Tree-Ring [wen2023tree]—or embedding messages into the noise distribution, as in Gaussian Shading [yang2024gaussian] and PRC-Watermark [gunn2024undetectable].
However, directly adapting image watermarking techniques to the video domain presents unique challenges. First, video DDIM inversion yields lower accuracy than image-based methods. And as a result, methods like VideoShield [hu2025videoshield] repeat watermark patterns in initial noise to enhance detection, but these compromise video quality and watermark invisibility. Second, watermark robustness suffers against temporal attacks like frame deletion or reordering, because treating videos as a single entity fails to localize watermarks temporally. Third, variable video lengths pose difficulties for algorithms relying on fixed noise initialization, limiting scalability.
@@
With these needs in mind, we introduce VideoMark, a distortion-free robust watermarking framework designed for video diffusion models. To achieve an imperceptible watermark that preserves the original noise characteristics, VideoMark utilizes pseudorandom error correction (PRC) codes [christ2024pseudorandom]. These codes map the watermark bits directly onto the initialized Gaussian noise for every frame. This specific design ensures the watermark integrates seamlessly, thus fulfilling our first design goal. To enable frame-specific watermarking, VideoMark processes each frame’s watermark independently while preserving sequential consistency across frames. Specifically, we generate an extended watermark message sequence. For each video, a random starting position within this master sequence initializes the first frame’s watermark, and subsequent frames derive their watermarks sequentially. This aligns with our second design objective, facilitating both individualized frame watermarking and temporal coherence.
To accurately extract the watermark, we propose a temporal matching module (TMM), which uses edit distance to align the decoded message with the embedded watermark sequence, thereby improving decoding accuracy. Even under temporal attacks such as frame deletion, TMM preserves the robustness of the embedded watermark.
In our experiments, we evaluate the effectiveness of our watermarking framework across different video diffusion models, demonstrating high decoding accuracy, high-quality generated videos, and strong invisibility. Our watermark achieves higher decoding accuracy compared to VideoShield, which is currently the state-of-the-art watermarking approach for video diffusion models. Additionally, our watermark achieves the best video quality on both the objective video evaluation benchmark VBench[huang2024vbench] and subjective assessments, maintaining parity with watermark-free videos. Importantly, our watermark remains undetectable to attackers without the key, ensuring stronger imperceptibility than other watermarking frameworks.
In summary, the contributions of this work are summerized as follows:
@@
2 Related Work
2.1 Video Diffusion Models
Diffusion models [sohl2015deep] progressively add noise to map data distributions to a Gaussian prior and recover original data via iterative denoising. Video diffusion models [ho2022video] use a 3D U-Net with interleaved spatial and temporal attention to generate high-quality, temporally consistent frames. Building on latent diffusion models [rombach2022high], SVD [blattmann2023stable] learns a multi-dimensional latent space for high-resolution frame synthesis. During generation, DDIM sampling [song2021denoising] efficiently reduces sampling steps while maintaining video quality compared to DDPM sampling [ho2020denoising].
Video diffusion models primarily follow two paradigms: Text-to-Video[wang2023modelscope, hu2025videoshield, huang2024vbench], where videos are generated based on text prompts, and Image-to-Video[2023i2vgenxl, hu2025videoshield, blattmann2023stable], where a video is generated starting from a single image. These paradigms enable the generation of realistic videos, but they also raise concerns regarding the potential generation of misleading content or copyright infringement.
2.2 Video Watermark
Video watermarking technology embeds imperceptible patterns into visual content and employs specialized detection methods to verify watermark presence [liu2024survey]. These methods are typically classified into two paradigms: post-processing and in-processing schemes.
Post-processing schemes introduce minimal visual perturbations, typically at the pixel level. Recent works [videoseal, luo2023dvmark, zhang2019robust, revmark] focus on training watermark embedding networks by optimizing discrepancies between watermarked and original videos, as well as encoding-decoding differences. However, these methods may struggle to balance trade-offs between video quality and watermark robustness.
In contrast to post-processing methods, in-processing schemes integrate the watermarking into the video generation process of current generative video models to better utilize their capabilities. For instance, Videoshield[hu2025videoshield] extends the Gaussian Shading technique[yang2024gaussian] from the image domain to the video domain, achieving improved robustness. However, repeating watermark bits during initialization induces fixed latent patterns, consequently degrading the quality of generated videos. Currently, only one in-processing video watermarking method exists, and prior approaches struggle to balance watermark robustness with video quality. We propose an undetectable video watermarking scheme to address this trade-off.
@@ 3.1 Diffusion Models
The reverse process is learned to gradually denoise from $\mathbf{x}T\sim\mathcal{N}(0,\mathbf{I})$ to generate content. While DDPM [ho2020denoising] introduces stochasticity in each denoising step, DDIM [song2020denoising] provides an approximately invertible deterministic sampling process:
$\begin{split}\mathbf{x}_{t-1}=\ &\sqrt{\alpha_{t-1}}\left(\frac{\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\,\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\alpha_{t}}}\right)\\ &+\sqrt{1-\alpha_{t-1}}\,\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\end{split}$ (2)
@@ 3.2 Pseudorandom Codes (PRC)
A PRC is a coding scheme that maps messages to statistically random-looking codewords. We adopt the construction from christ2024pseudorandom, which provides security based on the hardness of the Learning Parity with Noise (LPN) problem.
@@
Our implementation supports soft decisions on recovered bits, optimized for robust watermarking (see supplementary materials for details).
@@ 3.3 Generative Video Watermarking and Threat Model
2. Decoding: $\hat{m}=\mathcal{D}_{PRC}(V)$, where $\mathcal{D}_{PRC}$ extracts the watermark message $\hat{m}$ from the given video $V$.
[…]
@@ 4 Proposed Method
Figure 2: The overall framework of VideoMark. During the watermark embedding phase, $\epsilon$ denotes the standard Gaussian noise sampled randomly. In the I2V task, the first video frame prompts the prediction of initial noise during watermark extraction.
@@ 4.1 Watermark Generation
1:PRC-key $k$, number of frames $f$, channels $c$, height $h$, width $w$, message $m$, diffusion model $\mathcal{M}$, VAE decoder $\mathcal{D}$
2:Watermarked video $V$
@@
5:for $i=1$ to $f$ do
6:  Extract frame message $m_{i}$ from $M$ starting at position $p+i$
7:  Encode $m_{i}$ into PRC codeword $c_{i}\in\{-1,1\}^{c\times h\times w}$
@@
9:  Compute $\hat{\epsilon}_{i}\leftarrow c_{i}\cdot|\epsilon_{i}|$
@@
11:Denoise $\hat{\epsilon}$ using diffusion model $\mathcal{M}$ and decode with VAE decoder $\mathcal{D}$ to obtain video $V$
12:return $V$
In this section, we introduce the watermark generation process. VideoMark achieves high invisibility and video quality in diffusion-based watermarking by initializing each frame with pseudo-random Gaussian noise using PRC, followed by DDIM denoising[song2020denoising] and VAE decoding [vae]. To enhance diversity and adapt to varying video lengths, we employ an extended message list with a random start index.
Prior watermarking methods (e.g. VideoShield [hu2025videoshield]) often repeat identical noise patterns, compromising pseudo-randomness, reducing watermark bit capacity, and degrading invisibility and video quality. VideoMark addresses this by generating frame-specific pseudo-random initializations. For a video with $f$ frames, dimensions $c\times h\times w$ (channels, height, width), and a message bit $m_{i}^{\prime}$ per frame, the process is as follows. For each frame $i\in\{1,\ldots,f\}$, we sample Gaussian noise $\boldsymbol{\epsilon}_{i}\sim\mathcal{N}(0,\mathbf{I})\in\mathbb{R}^{c\times h\times w}$. Using a PRC key $k$, we encode $m_{i}^{\prime}$ to obtain a codeword $\mathbf{c}_{i}=\text{Encode}(k,m_{i}^{\prime})\in\{-1,1\}^{c\times h\times w}$. The watermarked noise is:
@@
$\begin{split}\hat{\boldsymbol{z}}_{i}^{(t-1)}=\ &\sqrt{\alpha_{t-1}}\left(\frac{\hat{\boldsymbol{z}}_{i}^{(t)}-\sqrt{1-\alpha_{t}}\,\boldsymbol{\epsilon}_{\theta}(\hat{\boldsymbol{z}}_{i}^{(t)},t)}{\sqrt{\alpha_{t}}}\right)\\ &+\sqrt{1-\alpha_{t-1}}\,\boldsymbol{\epsilon}_{\theta}(\hat{\boldsymbol{z}}_{i}^{(t)},t),\end{split}$ (3)
@@
To adapt to videos of varying lengths and increase diversity, we generate an extended message list $M=[m_{1},\ldots,m_{L}]$, where $L>f_{\text{max}}$ and $f_{\text{max}}$ is the maximum supported frame count. For each video, we sample a start index $p\sim\text{Uniform}(0,L-f)$, selecting messages $m_{i}^{\prime}=M[p+i]$ for $i\in\{1,\ldots,f\}$. These are encoded via PRC to produce $\mathbf{c}_{i}$. The random start index ensures diverse initializations across videos, improving security and reducing detectable patterns, while supporting arbitrary video lengths. This frame-wise approach resists temporal and spatial attacks. The pipeline is shown in Figure 2 and Algorithm 1.
[…]
Table 1: Main results of VideoMark. All columns present bit accuracy metrics except the Video Quality column.
@@ 4.2 Watermark Extraction
Decoding Function $\hat{m}=\mathcal{D}_{PRC}(V)$ extracts the embedded message from a watermarked video $V$ with $f$ frames. We first recover the approximate initial noise for each frame using the DDIM inverse process:
$\tilde{\boldsymbol{\epsilon}}_{i}=\mathcal{M}^{-1}(V_{i}),\quad i\in\{1,\ldots,f\}.$ (4)
@@
where the Decode function extracts the message bit encoded in the sign pattern using the PRC key $k$ (details of the PRC algorithm can be found in supplementary materials), matching the encoding process where $\hat{\boldsymbol{\epsilon}}_{i}=\mathbf{c}_{i}\cdot|\boldsymbol{\epsilon}_{i}|$ and $\mathbf{c}_{i}\in\{-1,1\}^{c\times h\times w}$. The complete decoded sequence is returned as $\hat{m}=[\hat{m}_{1},\ldots,\hat{m}_{f}]$.
[…]
@@
[…]
@@
$d_{N}(m_{i},\hat{m}_{j})=2\big(d_{H}(m_{i},\hat{m}_{j})-0.5\big)$ (8)
Since two random binary sequences are expected to have a Hamming distance of 0.5, this transformation linearly scales distances to the range [1,1][-1,1], enhancing the sensitivity of the detection mechanism.
[…]
$I_{j}=\arg\min_{j}\{d_{\text{edit}}(m,\hat{m}_{j})\},\quad j\in\{1,\ldots,f\}$ (9)
Then, using the indices, we identify both the starting index $s$ and the optimal alignment path between $m$ and $\hat{m}$:
[…]